# A solution to Banach conjecture

### v2

@@ Abstract.
In this paper, we first prove that the origin-symmetric star body with ellipsoidal sections is an ellipsoid. Then, we provide a complete proof of Banach’s isometric subspace problem in finite dimensions through the John ellipsoid.
@@ 1. Introduction
This problem goes back to Banach in 1932 [2, Remarks on Chapter XII] and is often referred to as the ”Banach conjecture.” The conjecture asserts that the answer is always affirmative. It has been confirmed in many, though not all, dimensions. Auerbach, Mazur, and Ulam established the conjecture for the case $n=2$ in [1]. Dvoretzky [5] later proved it for infinite-dimensional spaces $V$ and all $n\geq 2$. Gromov [7] showed that the answer is positive whenever $n$ is even or $\dim(V)\geq n+2$. Bor, Hernández-Lamoneda, Jiménez-Desantiago, and Montejano extended Gromov’s theorem to all $n$ congruent to $1$ modulo $4$, with the exception of $n=133$; see [3]. More recently, Ivanov, Mamaev, and Nordskova [8] obtained an affirmative solution for $n=3$. The problem can also be formulated for complex normed spaces; see [4], [7], [10] for further details.
@@
The proof of Theorem 1.1 is divided into two parts. Assume first that $K\cap\xi_{0}$ is an (n1)(n-1)-dimensional ellipsoid. Under this assumption, $K$ itself is an ellipsoid. Thus, to prove the main theorem, we choose a John ellipsoid in $K\cap\xi_{0}^{\perp}$ and then employ the linear invariance of the John ellipsoid to derive the final result.
@@ 2. Notation and Preliminaries
Let $\mathbb{R}^{n}$ be the classical $n$-dimensional Euclidean space. We denote by $\|x\|$ the Euclidean norm of $x\in\mathbb{R}^{n}$. We write $B^{n}=\{x\in\mathbb{R}^{n}:\|x\|\leq 1\}$ for the unit ball and $S^{n-1}=\{x\in\mathbb{R}^{n}:\|x\|=1\}$ for the unit sphere. We denote $\{e_{1},e_{2},\dots,e_{n}\}$ to be the standard basis in $\mathbb{R}^{n}$. For any $\xi\in S^{n-1}$, we denote $\xi^{\perp}:=\{x\in\mathbb{R}^{n}:\langle x,\xi\rangle=0\}$ to be a hyperplane that is perpendicular to $\xi$. Here, $\langle x,\xi\rangle$ is the usual inner product in $\mathbb{R}^{n}$. The Grassmann manifold of $k$-dimensional subspaces in $\mathbb{R}^{n}$ is denoted by $G(n,k)$. $GL(n)$ and $SL(n)$ are the general linear group and the special linear group in $\mathbb{R}^{n}$. For $\phi\in GL(n)$, we denote $|\phi|=|\det(\phi)|$ by the norm of the operator $\phi$.
@@ 3. Proof of the main theorem
Now we can prove the main theorem using John ellipsoids.
Theorem 3.2.
Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then the John ellipsoid $J(K\cap\xi_{0}^{\perp})$ can form an origin-symmetric star body $E$ by
$E:=\{\phi_{\xi}(J(K\cap\xi_{0}^{\perp})):\xi\in\mathbb{S}^{n-1}\}.$
Moreover, $E=J(K)$.
@@
First, we need to show the existence of the star body $E$. It is sufficient to show the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$.
@@
$\|\phi_{\xi}\|\leq\frac{\max_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}{\min_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}.$
Then we can find an arbitary convergent sequence $\{\xi_{n}\}_{n=1}^{\infty}$ with $\lim_{n\to\infty}\xi_{n}=\xi$ and pick $\phi_{\xi_{n}}\in\Phi(\xi_{n})$ such that there exists a subsequence $\{\xi_{n_{k}}\}_{k=1}^{\infty}$ with $\phi_{\xi_{n_{k}}}$ converging to $\phi\in GL(n)$ in operator norm. And thus, we have
@@
$d_{H}(\phi(K\cap\xi_{0}),K\cap\xi^{\perp})=0,$
which implies $\phi(K\cap\xi_{0})=K\cap\xi^{\perp}$, that is, $\phi\in\Phi(\xi)$. Therefore, $\Phi(\xi_{n_{k}})\to\Phi(\xi)$ as $k\to\infty$.
@@
Moreover, for any $g\in G_{K\cap\xi_{0}^{\perp}}$, $g(J(K\cap\xi_{0}^{\perp}))=J(g(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, where the John ellipsoid of $K\cap\xi_{0}^{\perp}$ is unique. Hence, $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$. Therefore, the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$ follows from the continuity of $\Phi(\xi)$.
Now, we need to show that for any $\xi,\eta\in\mathbb{S}^{n-1}$,
@@
holds for any $\theta\in\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\eta^{\perp}$. It is sufficient to show for any $\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ and any $\zeta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
@@
for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$.
Here we choose $\phi_{\zeta}\in\Phi(\zeta)$ with
@@
for some $g\in G_{K\cap\xi_{0}^{\perp}}$. The set $B$ consists of all $\xi$ for which there exists a vector $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ such that
@@
for every $g\in G_{K\cap\xi_{0}^{\perp}}$. It is clear that $A_{1}\cap A_{2}=\emptyset$. Moreover, the closedness of $A_{1}$ follows the closedness of $G_{K\cap\xi_{0}^{\perp}}$.
For $\xi\in A$ and $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
@@
for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$.
For $\xi\in A_{2}$ and $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
@@
where
$\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=\phi^{-1}_{\zeta}(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\cap\phi^{-1}_{\zeta}(\xi_{0}^{\perp}).$
Therefore, using the continuity of $\Phi$ and the fact that $\phi^{-1}_{\zeta}\neq g$, for any
$\eta=\frac{(1-t)\xi_{0}+t\eta_{\zeta}}{\|(1-t)\xi_{0}+t\eta_{\zeta}\|}\quad\text{with }t\in[0,1],\ \eta_{\zeta}\perp\phi^{-1}_{\zeta}(\xi_{0}^{\perp}),\ \text{and }\eta_{\zeta}\perp\xi_{0},$
there exists a linear map $\phi_{\eta}$ satisfying
$\phi_{\eta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\cap\eta^{\perp}$
and $\phi_{\eta}(\xi^{\perp})=\xi_{0}^{\perp}$. Therefore, we obtain
$\phi^{-1}_{\eta}(K\cap\xi_{0}^{\perp}\cap\eta^{\perp})=K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=K\cap\xi_{0}^{\perp}\cap\phi^{-1}_{\eta}(\xi_{0}^{\perp}),$
and thus, $\eta\in B$. Since $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ is compact, we can choose $\zeta_{0}$ such that $\|\eta_{\zeta_{0}},\xi_{0}\|$ reaches the maximum; then $\eta=\frac{(1-t)\xi_{0}+t\eta_{\zeta_{0}}}{\|(1-t)\xi_{0}+t\eta_{\zeta_{0}}\|}$ is in $A_{2}$.
Therefore, for any open set $O\subseteq A_{2}$ and any $\xi,\eta\in O$, there exists a linear isomorphism between $K\cap\xi_{0}^{\perp}\cap\xi^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap\eta^{\perp}$. If this were not the case, then by the continuity of $\Phi$ we could find some $\zeta\in O$ such that
$\phi_{\alpha}\bigl(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}\bigr)=g\bigl(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}\bigr)$
for every $\alpha\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi_{0},\zeta)$, where $g\in G_{K\cap\xi_{0}^{\perp}}$. Fixing $\xi\in O$, consider the set
$\{\eta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:K\cap\xi_{0}^{\perp}\cap\eta^{\perp}=\phi_{\eta}(K\cap\xi_{0}^{\perp}\cap\xi^{\perp}),\,\phi\in GL(n)\},$
which is closed. Consequently, for all $\xi,\eta\in\bar{O}$ there is a linear isomorphism between $K\cap\xi_{0}^{\perp}\cap\xi^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap\eta^{\perp}$. Then $\bar{O}\subseteq A_{2}$, which shows that $A_{2}$ is closed.
By the connectedness of $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we obtain either $A_{1}$ or $A_{2}$ as $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$. If $A_{1}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, the origin-symmetric star body
$E:=\{\phi_{\xi}(J(K\cap\xi_{0}^{\perp})):\xi\in\mathbb{S}^{n-1}\}$
exsits. Then, by Theorem 3.1, $E$ is an ellipsoid in $K$ with $E\cap\xi^{\perp}$ being the John ellipsoid of $K\cap\xi^{\perp}$ for any $\xi\in\mathbb{S}^{n-1}$.
@@
If $A_{2}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, then any (n2)(n-2)-dimensional sections of $K\cap\xi_{0}^{\perp}$ are mutually linearly dependent; then, using the linear dependence of all (n1)(n-1)-dimensional sections, we deduce that all (n2)(n-2)-dimensional sections of $K$ are likewise linearly dependent; therefore, by Gromov’s theorem [7], $K$ must be an ellipsoid. And hence, $E=K$ is an ellipsoid.
@@
For any ellipsoid $\tilde{E}$ contained in $K$ and any $\xi\in\mathbb{S}^{n-1}$, we have $|\tilde{E}\cap\xi^{\perp}|\leq|E\cap\xi^{\perp}|$, since $E\cap\xi^{\perp}$ serves as the John ellipsoid of $K\cap\xi^{\perp}$. As a result of the ellipsoid being an intersection body, Koldobsky and Zhang’s result (see [9] and [14]) yields $|\tilde{E}|\leq|E|$, and hence $E$ must be the John ellipsoid of $K$. ∎
Remark 3.1.

### v3

@@ Abstract.
In this paper, we begin by constructing an ellipsoid derived from the John ellipsoid of the sections. We then use this framework to present a full proof of Banach’s isometric subspace problem in finite-dimensional spaces via the John ellipsoid.
@@ 1. Introduction
This problem goes back to Banach in 1932 [2, Remarks on Chapter XII] and is often referred to as the ”Banach conjecture.” The conjecture asserts that the answer is always affirmative. It has been confirmed in many, though not all, dimensions. Auerbach, Mazur, and Ulam established the conjecture for the case $n=2$ in [1]. Dvoretzky [5] later proved it for infinite-dimensional spaces $V$ and all $n\geq 2$. Gromov [7] showed that the answer is positive whenever $n$ is even or $\dim(V)\geq n+2$. Bor, Hernández-Lamoneda, Jiménez-Desantiago, and Montejano extended Gromov’s theorem to all $n$ congruent to $1$ modulo $4$, with the exception of $n=133$; see [3]. More recently, Ivanov, Mamaev, and Nordskova [8] obtained an affirmative solution for $n=3$. The problem can also be formulated for complex normed spaces; see [4], [7], and [10] for further details.
@@
The proof of Theorem 1.1 is divided into two steps. We first select a John ellipsoid in $K\cap\xi_{0}^{\perp}$ and then use the linear invariance property of John ellipsoids to obtain a corresponding ellipsoid inside $K$. With this large ellipsoid at hand, we then establish the desired conclusion.
@@ 2. Notation and Preliminaries
Let $\mathbb{R}^{n}$ be the classical $n$-dimensional Euclidean space. We denote by $\|x\|$ the Euclidean norm of $x\in\mathbb{R}^{n}$. We write $B^{n}=\{x\in\mathbb{R}^{n}:\|x\|\leq 1\}$ for the unit ball and $S^{n-1}=\{x\in\mathbb{R}^{n}:\|x\|=1\}$ for the unit sphere. We denote $\{e_{1},e_{2},\dots,e_{n}\}$ to be the standard basis in $\mathbb{R}^{n}$. For any $\xi\in S^{n-1}$, we denote $\xi^{\perp}:=\{x\in\mathbb{R}^{n}:\langle x,\xi\rangle=0\}$ to be a hyperplane that is perpendicular to $\xi$. Here, $\langle x,\xi\rangle$ is the usual inner product in $\mathbb{R}^{n}$. The Grassmann manifold of $k$-dimensional subspaces in $\mathbb{R}^{n}$ is denoted by $G(n,k)$. $GL(n)$ and $SL(n)$ are the general linear group and the special linear group in $\mathbb{R}^{n}$. For $\phi\in GL(n)$, we denote $\|\phi\|_{op}=\max_{x\in B^{n}}\|\phi(x)\|$ by the operator norm of $\phi$.
@@ 3. Proof of the main theorem
Now we can prove the main theorem using John ellipsoids. First, we should introduce a couple of Lemmas.
Lemma 3.1.
Let $K\subseteq\mathbb{R}^{n}$ and $\phi\in GL(n)$. Then, for every $x\in\mathbb{R}^{n}$ we have
$\|x\|_{K}=\|\phi(x)\|_{\phi(K)}.$
@@
$\|\phi(x)\|_{\phi(K)}$ $=\inf\{\lambda:\phi(x)\in\lambda\phi(K)\}$
$=\inf\{\lambda:\phi(x)\in\phi(\lambda K)\}$
$=\inf\{\lambda:x\in\lambda K\}$
$=\|x\|_{K}.$
Lemma 3.2.
Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then the John ellipsoid $J(K\cap\xi_{0}^{\perp})$ can form an origin-symmetric star body $E$ by
$E:=\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}(J(K\cap\xi_{0}^{\perp})).$
Proof.
By the definition of $E$, the set $E$ is an origin-symmetric star set whose radial function is given by
$\rho_{E}(\theta)=\max\{\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta):\xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\}.$
Therefore, to prove that $\rho_{E}(\theta)$ is continuous, it suffices to establish the continuity of the mapping $\xi\mapsto\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$.
@@
$\|\phi_{\xi}\|_{op}\leq\frac{\max_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}{\min_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}.$
We may choose an arbitrary sequence $\{\xi_{n}\}_{n=1}^{\infty}$ converging to $\xi$, i.e., with $\lim_{n\to\infty}\xi_{n}=\xi$, and for each $n$ select $\phi_{\xi_{n}}\in\Phi(\xi_{n})$. Then there exists a subsequence $\{\xi_{n_{k}}\}_{k=1}^{\infty}$ such that the corresponding operators $\phi_{\xi_{n_{k}}}$ converge in operator norm to some $\phi\in GL(n)$. And thus, we have
@@
$d_{H}(\phi(K\cap\xi_{0}),K\cap\xi^{\perp})=0.$
This implies that $\phi(K\cap\xi_{0})=K\cap\xi^{\perp}$, and hence $\phi\in\Phi(\xi)$. Consequently, we have $\Phi(\xi_{n_{k}})\to\Phi(\xi)$ as $k\to\infty$.
@@
Moreover, for any $g\in G_{K\cap\xi_{0}^{\perp}}$, we have
$g(J(K\cap\xi_{0}^{\perp}))=J(g(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp}),$
since the John ellipsoid of $K\cap\xi_{0}^{\perp}$ is unique. Consequently, $J(K\cap\xi_{0}^{\perp})$ is invariant under the group $G_{K\cap\xi_{0}^{\perp}}$. Thus, the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ as a function of $\xi$ follows from the continuity of $\Phi(\xi)$. ∎
Remark 3.1.
Now we consider the origin-symmetric star body $E$. The next step is to confirm that $E\cap\xi^{\perp}$ indeed recovers $\phi_{\xi}(J(E\cap\xi_{0}^{\perp}))$.
Lemma 3.3.
Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then one of the following conditions must be satisfied.
  1. (1)
    For all $\xi,\eta\in\mathbb{S}^{n-1}$ we have
@@
    for every $\theta\in\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\eta^{\perp}$.
  2. (2)
    The body $K$ is an ellipsoid.
Proof.
Condition (1) can be reformulated as follows: for any $\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ and any $\zeta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we require that
@@
for every $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$, since, by choosing $\eta$ to be a basic direction, we have
$\phi_{\xi}(K\cap\xi^{\perp})=\phi_{\xi}\phi^{-1}_{\eta}(K\cap\eta^{\perp})$
for every $\xi\in\mathbb{S}^{n-1}$.
First, we introduce a function $\Gamma:\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\to\mathbb{R}_{\geq 0}$ defined by
$\Gamma(\xi)=\sup\Big\{\Big|\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}-\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}\Big|:\ \zeta,\eta\in\mathbb{S}^{n-1}\cap\,\mathrm{span}(\xi,\xi_{0})\Big\}.$
Then $\Gamma(\xi)$ is continuous because $\Phi(\xi)$ is continuous and $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$. Furthermore, $\Gamma(\xi)=\Gamma(\eta)$ whenever there exists a linear transformation between $K\cap\xi_{0}^{\perp}\cap\xi$ and $K\cap\xi_{0}^{\perp}\cap\eta$, as implied by Lemma 3.1.
Now fix $\xi$. For any $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$, select $\phi_{\zeta}\in\Phi(\zeta)$ such that
@@
for some $g\in G_{K\cap\xi_{0}^{\perp}}$. The set $A_{2}$ consists of all $\xi$ for which there exists a vector $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ such that
@@
for every $g\in G_{K\cap\xi_{0}^{\perp}}$. It is clear that $A_{1}\cap A_{2}=\emptyset$.
We once more partition $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ into two subsets, $B_{1}$ and $B_{2}$. Define
$B_{1}:=\{\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:\Gamma(\xi)=0\},\quad B_{2}:=\{\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:\Gamma(\xi)>0\}.$
Our goal is to show that $B_{1}$ and $B_{2}$ are closed.
Note that, for $\xi\in A_{1}$ and $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
@@
for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$. And thus, $\Gamma(\xi)=0$, so $A_{1}\subseteq B_{1}$, which implies $B_{2}\subseteq A_{2}$.
We now focus on the set $B_{2}$. For any $\xi\in B_{2}$, and therefore $\xi\in A_{2}$, and for any $\zeta,\eta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$ and $\zeta\neq\eta$, we have
@@
and, in addition,
$\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=\phi^{-1}_{\zeta}(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp})=\phi^{-1}_{\zeta}(K\cap\zeta^{\perp}\cap\eta^{\perp})=K\cap\xi_{0}^{\perp}\cap\phi^{-1}_{\zeta}(\eta^{\perp}).$
Hence, by exploiting the linearity of $\phi^{-1}_{\zeta}$, for any $y\in\phi^{-1}_{\zeta}(\eta^{\perp})$, we obtain $\langle\phi_{\zeta}(y),\eta\rangle=0$. Equivalently, this can be written as $y^{T}A^{T}_{\phi_{\zeta}}\eta=0$, which implies that $y\in(A^{T}_{\phi_{\zeta}}\eta)^{\perp}$, where $A_{\phi_{\zeta}}$ denotes the matrix associated with the linear transformation $\phi_{\zeta}$. Therefore, there exists a linear transform between $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap(A^{T}_{\phi_{\zeta}}\eta)^{\perp}$. Furthermore, the set
$E_{\zeta,\xi}:=\{A^{T}_{\phi_{\zeta}}\eta:\eta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})\}$
forms an ellipse, and the set
$C_{\zeta,\xi}:=\left\{\frac{A^{T}_{\phi_{\zeta}}\eta}{|A^{T}_{\phi_{\zeta}}\eta|}:\eta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})\right\}$
gives a circle. And thus, there exists a linear transform between $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap\tilde{\eta}^{\perp}$ for any $\tilde{\eta}\in C_{\zeta,\xi}$. Finally, we have
[…]
for any $\tilde{\eta}\in C_{\zeta,\xi}$.
Note that for any other point $\tilde{\xi}\in B_{2}$, there is a circle $C_{\tilde{\zeta},\tilde{\xi}}$ such that
$\Gamma(\tilde{\xi})=\Gamma(\bar{\eta})$
for all $\bar{\eta}\in C_{\tilde{\zeta},\tilde{\xi}}$. Consequently, the function $\Gamma$ must be constant on $B_{2}$, since any two circles on $\mathbb{S}^{n-1}$ intersect. Therefore, $B_{2}$ is closed.
By the connectedness of $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we obtain either $B_{1}$ or $B_{2}$ as $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$. If $B_{1}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we have $\Gamma(\xi)=0$ for any $\xi$, and hence Condition (1) holds.
@@
If $B_{2}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ holds, then $A_{2}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ implies that any (n2)(n-2)-dimensional sections of $K\cap\xi_{0}^{\perp}$ are all linearly dependent, because for every $\tilde{\eta}\in C_{\zeta,\xi}$ there is a linear transformation mapping $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}$ onto $K\cap\xi_{0}^{\perp}\cap\tilde{\eta}^{\perp}$. Using the fact that all (n1)(n-1)-dimensional sections are linearly dependent, we then conclude that every (n2)(n-2)-dimensional section of $K$ is also linearly dependent. Consequently, by Gromov’s theorem [7], $K$ must be an ellipsoid. ∎
To establish that $E\cap\xi^{\perp}=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$, it suffices to prove that $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ and $\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))$ coincide for any $k$-dimensional subspace ($1\leq k\leq n-2$). This leads to the following lemma, which is proved by induction.
Lemma 3.4.
Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then one of the following conditions must be satisfied.
  1. (1)
    For all $H\in G(n,k)$ with $1\leq k\leq n-2$, we have
    for every $\theta\in\mathbb{S}^{n-1}\cap H$ and $\xi,\eta\in H^{\perp}$.
  2. (2)
    The body $K$ is an ellipsoid.
Proof.
If $E$ is not an ellipsoid, then by Lemma 3.3, Condition (1) is satisfied for k=n2k=n-2. Now assume that Condition (1) holds for $k$.
For any $H\in G(n,k-1)$ and any $\xi\in H^{\perp}$, we have
$\mathrm{span}(H,\xi)\in G(n,k).$
Therefore, by the assumption of Condition (1), we obtain
for every $\theta\in\mathbb{S}^{n-1}\cap\,\mathrm{span}(H,\xi)$ and every $\zeta,\eta\in\bigl(\mathrm{span}(H,\xi)\bigr)^{\perp}$. In particular, this implies that
for all $\theta\in\mathbb{S}^{n-1}\cap H$ and all $\zeta,\eta\in\bigl(\mathrm{span}(H,\xi)\bigr)^{\perp}$, where
$\dim\bigl((\mathrm{span}(H,\xi))^{\perp}\bigr)=n-1-\dim(H)\geq 2.$
That Condition (1) is valid for k1k-1 follows from the fact that
$(\mathrm{span}(H,\xi_{1}))^{\perp}\cap(\mathrm{span}(H,\xi_{2}))^{\perp}\neq\emptyset$
for any $\xi_{1},\xi_{2}\in H^{\perp}$.
In particular, in the case $k=1$, for any $\theta\in\mathbb{S}^{n-1}$ and any $\xi,\eta\in\mathbb{S}^{n-1}\cap\theta^{\perp}$, we have
$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{E}(\theta).$
Consequently,
$E\cap\xi^{\perp}=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi^{\perp}).$
Now we can show the following theorem.
Theorem 3.2.
Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then the John ellipsoid $J(K\cap\xi_{0}^{\perp})$ can form an origin-symmetric star body $E$ by
$E:=\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}(J(K\cap\xi_{0}^{\perp})).$
Moreover, $E\cap\xi^{\perp}=J(K\cap\xi^{\perp})$ for any $\xi\in\mathbb{S}^{n-1}$ and $E=J(K)$.
Proof.
By Lemmas 3.2 and 3.4, either $K$ itself is an ellipsoid with $E=K$, or $E$ is an origin-symmetric star body satisfying $E\cap\xi^{\perp}=J(K\cap\xi^{\perp})$. In the latter case, Theorem 3.1 implies that $E$ is an ellipsoid.
@@
Now take an arbitrary ellipsoid $\tilde{E}$ contained in $K$. For every $\xi\in\mathbb{S}^{n-1}$, we have the inequality $|\tilde{E}\cap\xi^{\perp}|\leq|E\cap\xi^{\perp}|$, because $E\cap\xi^{\perp}$ is the John ellipsoid of the section $K\cap\xi^{\perp}$. Since the ellipsoid is an intersection body, the theorem of Koldobsky and Zhang (see [9] and [14]) implies that $|\tilde{E}|\leq|E|$. Consequently, $E$ is indeed the John ellipsoid of $K$. ∎